OpenStudy (anonymous):

Find all values a for which there exists an ordered pair (a, b) satisfying the following system of equations: a + ab^2 = 40b, and a - ab^2 = -32b.

OpenStudy (kittiwitti1):

\[a+ab^{2}\text{ & }=40b\text{, }a-ab^{2}\text{ & }=-32b\]Is this what you're asking for?

OpenStudy (loser66):

\(a(1+b^2)=40b\\a(1-b^2)=-32b\\\dfrac{40b}{1+b^2}=\dfrac{-32b}{1-b^2}\) Now, solve for b and then plug back to solve for a. If you are correct, the answer will be (12,3) for (a,b)

OpenStudy (ryamorgan284):

https://mathway.com/